A body of density 9.0 grams per centimeter cube(g/cm3) appears to have a mass of 27.0 grams in a liquid of density 1.2 grams per centimeter cube(g/cm3), what is the volume of the solid?
1
Expert's answer
2017-01-11T12:39:13-0500
ρ=m/V V=m/ρ
V(solid)=27.0/9.0=3 cm3